Mathematics 243 – Algebraic Structures Definitions Quiz 8 – November 10, 2006

Your Name:

A) (5) What is the definition of a (mathematical!) group.

B) (5) Let H = {[1], [3], [9]} in Z_{26}. Is H a group under multiplication mod 26?